Prove that every non-trivial subgroup of a cyclic group has finite index. Hence prove that (Q, +) is not cyclic.


1
Expert's answer
2021-05-07T14:37:38-0400

Let, "G" be a finite group of order "pq" , where "p<q" are primes.

Recall that the number "t_q" of sylow "q" -subgroup of "G"

satisfies "\frac{t_q}{p}" and "t_q=1"


"if\space t_q\not=1"

Then, "t_q\geqslant q+1>p" which contradicts the "\frac{t_q}{p}."

Thus "t_q=1"

Then G has normal sylow q-subgroup Q

we have "\frac{t_q}{p}"

and "t_q=1"

The first condition implies that "t_p=1" or "t_p=q"

The latter case implies that "q=1" which is excluded by our assumption that "p=q-1"

Thus, "t_p=1\space and \space G" has a normal sylow p-subgroup p.